Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB HED SQS SXP Sample Environment PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF MigratedByGroup
SCS Photon Commissioning Team XRO DOC WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 324     Entry time: 26 Aug 2023, 02:07
Author: Felix Brausse 
Group: MID 
Subject: Shift summary 

Shift summary the 25th of Aug

  • Found laser and X-ray spatiotemporal overlap with Bi diffraction peak
  • Found a diffraction peak from STO nano particles
  • X-ray dose test is ongoing

  • EXP_UPP_T10 was stuck at the hardware limit but the encoder value shows different values.

  • Cryo Failure at 1:30 am; BKR called and predicted ~20 hrs of downtime

ELOG V3.1.4-7c3fd00